# Natural Counterfactuals With Necessary Backtracking

### v2

@@ G.1 Possibility of Gratuitous Changes
G.2 Issues with the Distance Measure
@@
Appendix H Another Type of Minimal Change: Minimal Change in Local Causal Mechanisms.
@@
H.1 Concretization of Natural Counterfactuals: An Example Methodology
A Method Based on Mechanism Distance. We plugin mechanism distance Eqn. 9 into FIO framework Eqn. 2. Below is the equation of optimization:
$\min_{{\bf u}_{an({\bf A})^{*}}}\sum_{j}w_{j}||F({\bf u}_{j})-F({\bf u}^{*}_{j})||_{1}$ (10)
@@ H.1 Concretization of Natural Counterfactuals: An Example Methodology
$*\displaystyle s.t.\quad\epsilon<F({\bf u}^{*}_{j})<1-\epsilon,\forall{\bf u}^{*}_{j}\in u_{an({\bf A})^{*}}$
@@
$\mathcal{L}=\sum_{j}w_{j}||F({\bf u}_{j})-F({\bf u}^{*}_{j})||_{1}$ (11)
$+w_{\epsilon}\sum_{j}\max((\epsilon-F({\bf u}^{*}_{j}),0)+\max(\epsilon+F({\bf u}^{*}_{j})-1,0))$
@@
In the next section, we use Eqn. 11 for feasible intervention optimization across multiple machine learning case studies, showing that mechanism distance is as effective as perception distance, as discussed in the main paper.
H.2 Case Study
H.2.1 MorphoMNIST
@@
Intersection between Ours and NB (NCO=1, NB=1) (NCO=1, NB=0) (NCO=0, NB=1) (NCO=0, NB=0)
@@
We follow the experimental settings of simulation experiments in Sec. 6.1, except for two differences. One is that we use two state-of-the-art deep learning models, namely V-SCM (Pawlowski et al., 2020) and H-SCM (Ribeiro et al., 2023), as the backbones to learn counterfactuals. They use normalizing flows to learn causal relationships among $x$’s parent nodes, e.g., $(t,i)$ in MorphoMNIST. Further, to learn $p(x|t,i)$, notice that V-SCM uses VAE (Kingma & Welling, 2014) and HVAE (Maaløe et al., 2019). Another difference is that, instead of estimating the outcome with MAE, we follow the same metric called counterfactual effectiveness in Ribeiro et al. (2023) developed by Monteiro et al. (2023), First, trained on the dataset, parent predictors given a value of $x$ can predict parent values, i.e., $(t,i)$’s, and then measure the absolute error between parent values after hard intervention or feasible intervention and their predicted values, which is measured on image the Learned SCM generates given the input of $(t,i)$.
@@
Figure 12: Causal Graph and samples of Morpho-MNIST.
@@
H.2.2 3DIdentBOX
Table 14: Results on Weak-3DIdent and Stong-3DIdent
Dataset Counterfactuals $d$ $h$ $v$ $\gamma$ $\alpha$ $\beta$ $b$
Ours 0.0241 0.0182 0.0339 0.348 0.224 0.0371 0.00416
Ours 0.0633 0.0512 0.0518 0.326 0.348 0.151 0.00464
In this task, we utilize practical public datasets called 3DIdentBOX, which encompass multiple datasets (Bizeul et al., 2023). Specifically, we employ Weak-3DIdent and Strong-3DIdent, both of which share the same causal graph depicted in Fig. 13 (a), consisting of an image variable denoted as $x$ and seven parent variables. These parent variables, denoted as $(d,h,v)$, control the depth, horizon position, and vertical position of the teapot in image $x$ respectively. Additionally, the variables $(\gamma,\alpha,\beta)$ govern three types of angles associated with the teapot within images, while variable $b$ represents the background color of the image. As illustrated in Fig. 13 (a), causal relationships exist among three pairs of parent variables, i.e., $(h,d)$, $(v,\beta)$ and $(\alpha,\gamma)$. It is important to note a distinction between Weak-3DIdent and Strong-3DIdent. In Weak-3DIdent, there exists a weak causal relationship between the variables of each pair, as shown in Fig. 13 (b), whereas in Strong-3DIdent, the causal relationship is stronger, as depicted in Fig. 13 (c).

### v3

@@ F.1 Possibility of Gratuitous Changes
F.2 Issues with the Distance Measure
@@
Appendix G Another Type of Minimal Change: Minimal Change in Local Causal Mechanisms.
@@
G.1 Concretization of Natural Counterfactuals: An Example Methodology
A Method Based on Mechanism Distance. We plugin mechanism distance Eqn. 11 into FIO framework Eqn. 2. Below is the equation of optimization:
$\min_{{\bf u}_{an({\bf A})^{*}}}\sum_{{\bf u}_{j}\in{\bf u}_{an({\bf A})},{\bf u}_{j}^{*}\in{\bf u}_{an({\bf A})^{*}}}w_{j}|F({\bf u}_{j})-F({\bf u}^{*}_{j})|$ (12)
@@ G.1 Concretization of Natural Counterfactuals: An Example Methodology
$s.t.\quad\epsilon<F({\bf u}^{*}_{j})<1-\epsilon,\forall{\bf u}^{*}_{j}\in u_{an({\bf A})^{*}}$
@@
$\mathcal{L}=\sum_{{\bf u}_{j}\in{\bf u}_{an({\bf A})},{\bf u}_{j}^{*}\in{\bf u}_{an({\bf A})^{*}}}w_{j}|F({\bf u}_{j})-F({\bf u}^{*}_{j})|$ (13)
$+w_{\epsilon}\sum_{{\bf u}_{j}\in{\bf u}_{an({\bf A})},{\bf u}_{j}^{*}\in{\bf u}_{an({\bf A})^{*}}}\max((\epsilon-F({\bf u}^{*}_{j}),0)+\max(\epsilon+F({\bf u}^{*}_{j})-1,0))$
@@
In the next section, we use Eqn. 13 for feasible intervention optimization across multiple machine learning case studies, showing that mechanism distance is as effective as perception distance, as discussed in the main paper.
G.2 Case Study
G.2.1 MorphoMNIST
@@
Intersection between Ours and NB (NC=1, NB=1) (NC=1, NB=0) (NC=0, NB=1) (NC=0, NB=0)
@@
We follow the experimental settings of simulation experiments in Sec. 6.1, except for two differences. One is that we use two state-of-the-art deep learning models, namely V-SCM (Pawlowski et al., 2020b) and H-SCM (Ribeiro et al., 2023), as the backbones to learn counterfactuals. They use normalizing flows to learn causal relationships among $x$’s parent nodes, e.g., $(t,i)$ in MorphoMNIST. Further, to learn $p(x|t,i)$, notice that V-SCM uses VAE (Kingma & Welling, 2014) and HVAE (Maaløe et al., 2019). Another difference is that, instead of estimating the outcome with MAE, we follow the same metric called counterfactual effectiveness in (Ribeiro et al., 2023) developed by (Monteiro et al., 2023), First, trained on the dataset, parent predictors given a value of $x$ can predict parent values, i.e., $(t,i)$’s, and then measure the absolute error between parent values after hard intervention or LBF intervention and their predicted values, which is measured on image the pretrained causal model generates given the input of $(t,i)$.
@@
Figure 8: Causal Graph and samples of MorphoMNIST.
@@
G.2.2 3DIdentBOX
Table 15: Results on Weak-3DIdent and Stong-3DIdent
Dataset Counterfactuals $d$ $h$ $v$ $\gamma$ $\alpha$ $\beta$ $b$
Ours 0.0241 0.0182 0.0339 0.348 0.224 0.0371 0.00416
Ours 0.0633 0.0512 0.0518 0.326 0.348 0.151 0.00464
In this task, we utilize practical public datasets called 3DIdentBOX, which encompass multiple datasets (Bizeul et al., 2023). Specifically, we employ Weak-3DIdent and Strong-3DIdent, both of which share the same causal graph depicted in Fig. 9 (b), consisting of an image variable denoted as $x$ and seven parent variables. These parent variables, denoted as $(d,h,v)$, control the depth, horizon position, and vertical position of the teapot in image $x$ respectively. Additionally, the variables $(\gamma,\alpha,\beta)$ govern three types of angles associated with the teapot within images, while variable $b$ represents the background color of the image. As illustrated in Fig. 9 (a), causal relationships exist among three pairs of parent variables, i.e., $(h,d)$, $(v,\beta)$ and $(\alpha,\gamma)$. It is important to note a distinction between Weak-3DIdent and Strong-3DIdent. In Weak-3DIdent, there exists a weak causal relationship between the variables of each pair, as shown in Fig. 9 (b), whereas in Strong-3DIdent, the causal relationship is stronger, as depicted in Fig. 9 (c).